# Knowledge Sharing and Transfer via Centralized Reward Agent for Multi-Task Reinforcement Learning

### v1

@@ 4.1.3 Policy Agents with Knowledge Rewards
$r^{pol}_{i}=r^{env}_{i}+\lambda r^{knw},\quad r^{knw}\sim\pi^{rwd}(\cdot|s_{i},a_{i}),$
@@
${\pi^{pol}_{i}}^{*}=\operatorname*{arg\,max}_{\pi^{pol}_{i}}\underset{(s_{i},a_{i})\sim\pi^{pol}_{i}}{\mathbb{E}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r^{pol}_{i}\Big{]}.$
@@ 4.2 Information Synchronization of Policy Agents
$s_{i}=\frac{\bm{c}^{T}\cdot\bm{H}_{i}}{\sqrt{D}},\quad\bm{c}=\frac{1}{N}\sum_{k=1}^{N}\bm{H}_{k},$
@@ 4.3 Overall Framework
9:     Get similarity and performance weights $\bm{w}^{sim}$, $\bm{w}^{per}$
10:     $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$
11:     $\{s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i}\}_{\mathcal{B}}\sim\mathcal{D}|\bm{w}$triangleright Draw a batch of samples based on the sampling weight
12:     Update centralized reward agent $\mathcal{A}^{rwd}$
13:  end for
@@ 5 Experiments
We conduct experiments in three environments with multiple tasks: 2DMaze and 3DPickup by Chevalier-Boisvert et al., (2024), and MujocoCar by Ji et al., (2023), which include both discrete and continuous control problems. All tasks provide challenging sparse environmental rewards, where the agent only receives a reward of $1$ upon successful completion of the final objective, and a reward of 00 is given in all other states. Detailed configurations are shown in Figure 3, with each task’s setting provided in Appendix A.
(a) 2DMaze
(b) 3DPickup
(c) MujocoCar
Figure 3: Experimental environments with multiple tasks. (a) 2DMaze: A $10\times 10$ 2D maze environment where the agent must pick up a key and then pass through a door to exit. Different tasks involve different placements of walls, keys, and doors. (b) 3DPickup: A 3D environment where the agent’s task is to navigate to and pick up target objects. Tasks vary by different object types, e.g., balls, cubes, keys and health kits, and their locations. (c) MujocoCar: A mujoco-based race car aims to navigate to specified areas, where each task has a different target location.
@@ 5.1 Comparative Evaluation in MTRL
In this section, we evaluate the learning performance of CenRA in the MTRL context, configuring each environment with four different tasks. We benchmark CenRA against several state-of-the-art baselines, including: (a) the backbone algorithms of the policy agents in CenRA: the Deep Q-Network (DQN) Mnih et al., (2015) for the discrete control tasks in 2DMaze and 3DPickup, and the Soft Actor-Critic (SAC) Haarnoja et al., 2018b for the continuous control tasks in MujocoCar; (b) The RL with an Assistant Reward Agent (ReLara) algorithm Ma et al., 2024b , which learns an additional reward agent to generate informative rewards. ReLara can be considered as a special case of CenRA where each policy agent is paired with a separate reward agent, without cross-task information sharing; (c) The Policy Optimization and Policy Correction (PiCor) algorithm Bai et al., (2023), which uses policy correction to adapt and transfer information across policies among individual tasks; (d) The Multi-Critic Actor Learning (MCAL) algorithm Mysore et al., (2022), which maintains separate critics for each task for value estimation and knowledge representation, while training a unified multi-task actor. We use the CleanRL library Huang et al., (2022) or the author-provided codes to implement these baselines.
(a) 2DMaze
(b) 3DPickup
(c) MujocoCar
@@
Figure 4: Comparison of the learning performance of CenRA with the baselines.
Table 1: Comparison of episodic returns and standard errors.
Algo. 2DMaze
Maze #1 Maze #2 Maze #3 Maze #4
CenRA 0.89 $\pm$ 0.03 0.91 $\pm$ 0.02 0.92 $\pm$ 0.02 0.93 $\pm$ 0.02
DQN 0.77 $\pm$ 0.05 0.49 $\pm$ 0.09 0.54 $\pm$ 0.09 0.78 $\pm$ 0.05
ReLara 0.83 $\pm$ 0.05 0.88 $\pm$ 0.03 0.76 $\pm$ 0.05 0.74 $\pm$ 0.05
PiCor 0.80 $\pm$ 0.05 0.85 $\pm$ 0.05 0.78 $\pm$ 0.06 0.84 $\pm$ 0.05
MCAL 0.76 $\pm$ 0.07 0.72 $\pm$ 0.09 0.57 $\pm$ 0.08 0.69 $\pm$ 0.08
Algo. 3DPickup
Ball Cube Key Health kit
CenRA 0.95 $\pm$ 0.02 0.68 $\pm$ 0.09 0.80 $\pm$ 0.06 0.69 $\pm$ 0.07
DQN 0.34 $\pm$ 0.06 0.07 $\pm$ 0.03 0.28 $\pm$ 0.03 0.28 $\pm$ 0.07
ReLara 0.45 $\pm$ 0.08 0.39 $\pm$ 0.10 0.33 $\pm$ 0.09 0.29 $\pm$ 0.08
PiCor 0.28 $\pm$ 0.07 0.44 $\pm$ 0.08 0.35 $\pm$ 0.09 0.28 $\pm$ 0.10
MCAL 0.24 $\pm$ 0.07 0.34 $\pm$ 0.05 0.36 $\pm$ 0.08 0.45 $\pm$ 0.07
Algo. MujocoCar
Target #1 Target #2 Target #3 Target #4
CenRA 588.2 $\pm$ 0.73 549.3 $\pm$ 0.64 447.7 $\pm$ 0.67 474.3 $\pm$ 0.66
SAC 247.3 $\pm$ 0.53 210.6 $\pm$ 0.49 177.7 $\pm$ 0.41 156.4 $\pm$ 0.38
ReLara 434.3 $\pm$ 0.68 436.0 $\pm$ 0.67 308.1 $\pm$ 0.58 420.8 $\pm$ 0.69
PiCor 411.0 $\pm$ 0.66 426.0 $\pm$ 0.63 470.7 $\pm$ 0.66 442.5 $\pm$ 0.70
MCAL 341.0 $\pm$ 0.62 315.6 $\pm$ 0.58 430.4 $\pm$ 0.59 389.8 $\pm$ 0.59
@@
We train multiple instances with different random seeds for each task to report the average data. The results of the episodic returns and their standard errors throughout training are shown in Figure 4 and Table 1. To ensure a fair comparison, we use the same hyperparameters (if applicable) and network structures across all experiments, the details are provided in Appendix B.
We observe that CenRA consistently outperforms the baselines in three main aspects. First, CenRA achieves the highest episodic returns in nearly all tasks, demonstrating superior learning efficiency and faster convergence. Importantly, it also maintains good stability and robustness, which can be seen in the significantly fewer fluctuations and oscillations especially after convergence, compared to other models. Second, unlike PiCor and MCAL, which have shown uneven progress across different tasks in the same environment, CenRA maintains a well-balanced performance by showing relatively consistent learning progress and minimal variability in learning outcomes across each four-task group. This ensures that no single task’s performance is disproportionately high at the expense of others, which is crucial in multi-task learning. Third, the centralized reward agent in CenRA effectively enhances knowledge sharing among tasks. This is evident from the comparison with ReLara, which uses independent reward agents and lacks the mechanism for knowledge exchange. By extracting and distributing insights from one task to another, the CRA improves the learning efficiency of individual tasks, highlighting the advantages of integrated knowledge management.
5.2 What Has the Centralized Reward Agent Learned?
@@ 5.3 Knowledge Transfer to New Tasks
In this section, we assess the centralized reward agent’s ability to transfer learned knowledge to new tasks. We continue using the CRA model trained on the four tasks in Section 5.1 but initiate a new policy agent to tackle a newly designed task that the models have never encountered before. For the CenRA framework, we explore two scenarios: one where the CRA continues to optimize while cooperating with the new policy agent (CenRA w/ learning), and another where only the policy agent is trained without further updates to the CRA (CenRA w/o learning). We compare the two approaches with the backbone algorithms (DQN and SAC) and ReLara. In the ReLara model, the reward agent starts training anew (without pre-learned knowledge). The results are shown in Figure 6 and Table 2.
We observe that CenRA with further learning achieved rapid convergence, mainly due to the CRA’s retention of previously acquired knowledge and its ability to quickly adapt to new tasks through continuous optimization. Notably, even without any additional training, CenRA outperforms both the ReLara, which requires training a new reward agent, and the backbone algorithms, which lack additional information. In the CenRA framework, the CRA stores environment-relevant knowledge and transfers it to other policy agents, thus improving their learning efficiency. Additionally, sharing pre-learned information with policy agents is particularly crucial in our experiments with challenging sparse rewards. Without any external knowledge, learning would require extensive exploration. However, the CRA provides meaningful dense rewards that significantly accelerate the learning process, even during the initial phases.
@@
Table 2: Comparison of the episodic returns and standard errors in new tasks.
Algo. New 2DMaze New 3DPickup New MujocoCar
CenRA w/ learning 0.95 $\pm$ 0.01 0.96 $\pm$ 0.00 532.00 $\pm$ 0.61
CenRA w/o learning 0.89 $\pm$ 0.03 0.68 $\pm$ 0.00 524.73 $\pm$ 0.60
ReLara 0.76 $\pm$ 0.06 0.26 $\pm$ 0.00 224.65 $\pm$ 0.49
DQN/SAC 0.26 $\pm$ 0.08 0.16 $\pm$ 0.00 129.06 $\pm$ 0.30
@@
To further demonstrate CenRA’s transferability, we select the 2DMaze environment to visualize the learned knowledge of CRA in an unseen task. As shown in Figure 7, we plot the directions of actions that yield the maximum knowledge reward at each position, categorized into two scenarios before and after obtaining the key. While some guidance in peripheral areas may be slightly off-target, most states receive reasonable rewards that align with human understanding. This demonstrates the effective transfer of knowledge and with such dense rewards, the agent’s adaptation to new tasks is able to be well supported.
@@ 5.4 Effect of Sampling Weight
We conducted experiments to understand the effects of the information synchronization mechanism in the CenRA framework. Specifically, we compare the full CenRA model against three variants: (a) CenRA without the similarity weight $\bm{w}^{sim}$, (b) CenRA without the performance weight $\bm{w}^{per}$, and (c) CenRA without the entire sampling weight. The comparison results are shown in Figure 8 and Table 3.
(a) 2DMaze
(b) 3DPickup
(c) MujocoCar
Figure 8: Comparison of CenRA with ablation of different batch sampling control weights.
@@
Both weights are essential for the information synchronization mechanism, with the performance weight $\bm{w}^{per}$ having a more important effect. This weight enables CRA to pay more attention to the policy agents that are underperforming or progressing slowly. This suggests that considering the overall learning performance of the multi-task system is also an important objective CenRA seeks to achieve.
@@ 6 Discussion and Conclusion
We propose a novel framework named CenRA that innovatively integrates reward shaping into MTRL. The framework not only can share environment-relevant knowledge across tasks to improve learning efficiency, but also effectively addresses the sparse-reward challenge. Specifically, a CRA functions as a knowledge pool, responsible for knowledge distillation and distribution across multiple tasks. Furthermore, our information synchronization mechanism mitigates knowledge distribution disparities across tasks, ensuring optimal system-wide performance. Experiments demonstrate that the dense knowledge rewards generated by the CRA efficiently guide the policy agents’ learning, leading to faster convergence compared to baseline methods. The CenRA framework proves to be robust and performs superiorly in both multi-task learning and new tasks adaptation.
CenRA’s current limitation is its requirement for uniform state and action spaces across tasks. Future improvements could explore preprocessing techniques to adapt the framework for varying state and action spaces to broaden CenRA’s applicability. Additionally, the fixed trade-off between similarity weight and performance weight may not be ideal. A more flexible approach, potentially through adaptive regulation of these weights, could enhance the framework further. Besides, the performance weight might favor underperforming tasks, achieving overall balance but potentially limiting the upper bound of top-performing tasks. This issue could necessitate a better mechanism for trade-off.
7 Acknowledgement
@@ Appendix A Mutli-Task Experimental Configurations
In Section 5 Experiments, we conduct experiments in three environments with multiple tasks: 2DMaze, 3DPickup, and MujocoCar. The detailed configurations of each task are illustrated in Figure 9.
Figure 9: The configurations of multiple tasks used in your experiments.
Appendix B Network Structures and Hyperparameters

### v2

@@ 4.1.3 Policy Agents with Knowledge Rewards
$r^{pol}_{i}=r^{env}_{i}+\lambda r^{knw},\quad r^{knw}\sim\pi^{rwd}(\cdot|s_{i},a_{i}),$ (3)
@@
${\pi^{pol}_{i}}^{*}=\operatorname*{arg\,max}_{\pi^{pol}_{i}}\underset{(s_{i},a_{i})\sim\pi^{pol}_{i}}{\mathbb{E}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r^{pol}_{i}\Big{]}.$ (4)
@@ 4.2 Information Synchronization of Policy Agents
$s_{i}=\frac{\bm{c}^{T}\cdot\bm{H}_{i}}{\sqrt{D}},\quad\bm{c}=\frac{1}{N}\sum_{k=1}^{N}\bm{H}_{k},$ (5)
@@ 4.3 Overall Framework
9:     $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$triangleright Calculate sampling weight
10:     $\{s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i}\}_{\mathcal{B}}\sim\mathcal{D}|\bm{w}$triangleright Draw samples based on the sampling weight
11:     Update centralized reward agent $\mathcal{A}^{rwd}$
12:  end for
@@ 5 Experiments
We conduct experiments in four MTRL domains as shown in Figure 3: the widely used Meta-World benchmark (including ML10 with 10 tasks and ML50 with 50 tasks) (Yu et al., 2020b, ), 2DMaze, 3DPickup (Chevalier-Boisvert et al.,, 2024), and MujocoCar (Ji et al.,, 2023). All tasks, including those in Meta-World, are crafted to provide sparse environmental rewards, where the agent receives a reward of $1$ only upon successful completion of the final objective, and 00 otherwise. The detailed task configurations are provided in Appendix A.
(a) Meta-World (ML10-sparse and ML50-sparse)
(b) 2DMaze
(c) 3DPickup
(d) MujocoCar
Figure 3: Environments with multiple tasks. (a) Meta-World: two sparse-reward versions are used: ML10-sparse and ML50-sparse, including diverse robotic manipulation tasks. (b) 2DMaze: 2D maze tasks where the agent must pick up a key and then pass through a door to exit. (b) 3DPickup: 3D maze tasks where the agent aims to navigate to and pick up different target objects at different locations. (c) MujocoCar: mujoco-based race car aims to navigate to different specified areas.
@@ 5.1 Comparative Evaluation in MTRL
We benchmark CenRA against several state-of-the-art baselines: (a) the backbone RL algorithms of the policy agents: DQN (Mnih et al.,, 2015) for discrete control tasks and SAC (Haarnoja et al., 2018b, ) for continuous control tasks; (b) the ReLara algorithm (Ma et al., 2024a, ), which can be regarded as a decentralized variant of CenRA, where each policy agent is paired with a separate reward agent, without cross-task information sharing; (c) the Policy Optimization and Policy Correction (PiCor) algorithm (Bai et al.,, 2023); (d) the Multi-Critic Actor Learning (MCAL) algorithm (Mysore et al.,, 2022); (e) the Parameter-compositional MTRL (PaCo) algorithm (Sun et al.,, 2022); and (f) the MTRL with Soft Modularization (SoftModule) (Yang et al.,, 2020). They are implemented by either the CleanRL library (Huang et al.,, 2022) or official codebases. Each task is trained with 10 different random seeds, and the average results are reported.
In the Meta-World domain, ML10-sparse provides 10 training tasks and 5 held-out test tasks, while ML50-sparse includes 45 training tasks and 5 test tasks. For the remaining domains, each consists of 4 training tasks and 1 test task. In this section, we evaluate the final returns achieved by the trained agents, averaged over all training tasks in each domain, as shown in Table 1. We additionally report the episodic returns and their standard errors throughout training in the 2DMaze, 3DPickup, and MujocoCar domains in Figure 4. To ensure a fair comparison, we adopt consistent hyperparameters (where applicable) and identical network architectures across all experiments; detailed configurations are provided in Appendix B.
@@
Figure 4: Comparison of CenRA with baselines in 2DMaze, 3DPickup, and MujocoCar domains.
Table 1: Episodic returns (mean $\pm$ standard error) of all trained agents tested over 100 episodes and averaged across all training tasks in each domain ($\uparrow$ higher is better).
Algorithm ML10-sparse ML50-sparse 2DMaze 3DPickup MujocoCar
CenRA (ours) 0.875 $\pm$ 0.121 0.755 $\pm$ 0.034 0.913 $\pm$ 0.023 0.880 $\pm$ 0.060 514.875 $\pm$ 0.675
DQN/SAC 0.256 $\pm$ 0.056 0.189 $\pm$ 0.012 0.645 $\pm$ 0.070 0.243 $\pm$ 0.048 198.000 $\pm$ 0.453
ReLara 0.674 $\pm$ 0.105 0.541 $\pm$ 0.057 0.803 $\pm$ 0.065 0.565 $\pm$ 0.088 429.800 $\pm$ 0.655
PiCor 0.865 $\pm$ 0.230 0.672 $\pm$ 0.123 0.818 $\pm$ 0.053 0.438 $\pm$ 0.085 437.550 $\pm$ 0.663
MCAL 0.842 $\pm$ 0.067 0.605 $\pm$ 0.055 0.885 $\pm$ 0.080 0.548 $\pm$ 0.068 369.200 $\pm$ 0.595
PaCo 0.854 $\pm$ 0.045 0.582 $\pm$ 0.022 0.834 $\pm$ 0.057 0.557 $\pm$ 0.072 421.210 $\pm$ 0.635
SoftModule 0.630 $\pm$ 0.042 0.423 $\pm$ 0.057 0.822 $\pm$ 0.076 0.486 $\pm$ 0.055 355.125 $\pm$ 0.594
We observe that CenRA consistently outperforms all baselines in three main aspects. First, it achieves the highest episodic returns in all tasks, demonstrating superior learning efficiency and faster convergence. Moreover, it demonstrates good stability and robustness, exhibiting fewer fluctuations and oscillations, especially after convergence, compared to other models. Notably, all tasks provide only sparse rewards, CenRA addresses this challenge through the auxiliary dense rewards with meaningful information, effectively guiding learning. This mechanism not only distinguishes CenRA from other structurally shared methods, but also provides a targeted solution to the sparse-reward problem. Second, while baselines like PiCor and MCAL often show uneven progress across different tasks within the same domain, CenRA maintains well-balanced performance by showing relatively consistent learning progress and minimal variability across each four-task groups. This ensures that no single task dominates or lags behind, which is crucial in multi-task learning. Third, the CRA effectively enhances knowledge sharing among tasks. This is evident from the comparison with ReLara, which uses independent reward agents and lacks the mechanism for knowledge exchange. By extracting and distributing insights from one task to another, the CRA improves the learning efficiency of individual tasks, highlighting the advantages of integrated knowledge management.
5.2 What Has the Centralized Reward Agent Learned?
@@ 5.3 Knowledge Transfer to New Tasks
In this section, we assess the CRA’s ability to transfer previously learned knowledge to unseen tasks. Specifically, we continue using the trained CRA model in Section 5.1, while initializing new policy agents to tackle new tasks from the same domain. These include 5 test tasks for ML10-sparse and ML50-sparse, and 1 test task for each of the remaining domains, none of which were encountered during the initial training. For the CenRA, we explore two scenarios: (1) the CRA continues to be optimized in collaboration with the new policy agent (CenRA w/ learning); and (2) only the policy agent is updated while the CRA remains fixed, relying only on its previously acquired knowledge (CenRA w/o learning). We compare the two settings against the backbone algorithms and ReLara. In ReLara, the reward agent is trained anew without pre-learned knowledge. The results are presented in Figure 6 and Table 2.
We observe that CenRA with further learning achieved rapid convergence, mainly due to the CRA’s ability to retain previously acquired knowledge while continuing to adapt to new tasks through ongoing optimization. Remarkably, even without any additional training, CenRA still outperforms both ReLara, which requires training a new reward agent, and the backbone algorithms, which lack additional information. This advantage stems from the CRA’s ability to encode and transfer environment-relevant knowledge, which can then be directly reused by new policy agents to guide their learning. Such knowledge transfer is particularly critical in our experiments involving challenging sparse-reward tasks. Without any external knowledge, learning would require extensive exploration. However, the CRA provides meaningful dense rewards that significantly accelerate the learning process, even during the initial phases.
@@
Table 2: Episodic returns (mean $\pm$ standard error) of all trained agents in the new tasks, tested over 100 episodes in each domain ($\uparrow$ higher is better).
Algorithm ML10-sparse ML50-sparse 2DMaze 3DPickup MujocoCar
CenRA w/ learning 0.902 $\pm$ 0.021 0.824 $\pm$ 0.012 0.952 $\pm$ 0.010 0.963 $\pm$ 0.002 532.080 $\pm$ 1.610
CenRA w/o learning 0.887 $\pm$ 0.011 0.809 $\pm$ 0.009 0.894 $\pm$ 0.032 0.678 $\pm$ 0.003 524.727 $\pm$ 0.588
ReLara 0.702 $\pm$ 0.086 0.612 $\pm$ 0.012 0.759 $\pm$ 0.056 0.263 $\pm$ 0.002 224.648 $\pm$ 0.492
DQN/SAC 0.228 $\pm$ 0.105 0.210 $\pm$ 0.034 0.263 $\pm$ 0.084 0.158 $\pm$ 0.003 129.055 $\pm$ 0.296
@@
To further demonstrate CenRA’s transferability, we select the 2DMaze environment to visualize the knowledge provided by CRA when facing an unseen task. As shown in Figure 7, we plot the directions of actions that yield the maximum knowledge reward at each position, categorized into two scenarios: before and after obtaining the key. While some guidance in peripheral regions may appear slightly misaligned, most states receive reasonable rewards that align with human understanding. This demonstrates the effectiveness of knowledge transfer, and with such dense rewards, the agent’s adaptation to new tasks is able to be well supported.
@@ 5.4 Effect of Sampling Weight
We conducted experiments to understand the effects of the information synchronization mechanism in the CenRA. Specifically, we compare the full CenRA model against three variants: (a) CenRA without the similarity weight $\bm{w}^{sim}$, (b) CenRA without the performance weight $\bm{w}^{per}$, and (c) CenRA without the entire sampling weight. To better illustrate the differences among tasks and highlight the role of sampling weights in task coordination and synchronization, we select the four-task domains, i.e., 2DMaze, 3DPickup, and MujocoCar. The comparison results are shown in Table 3.
@@
Both weights play essential roles in the information synchronization, with the performance weight $\bm{w}^{per}$ having a more important impact. It allows CRA to pay more attention to the policy agents that are underperforming or progressing slowly. This suggests that considering the overall learning performance of the multi-task system is also an important objective that CenRA seeks to achieve.
@@ 6 Discussion and Conclusion
We propose a novel framework named CenRA that innovatively integrates reward shaping into MTRL. The framework not only shares domain knowledge across tasks to improve learning efficiency but also effectively addresses the sparse-reward challenge. Specifically, the centralized reward agent functions as a knowledge pool, responsible for knowledge distillation and distribution across multiple tasks. Furthermore, our information synchronization mechanism mitigates knowledge distribution disparities across tasks, ensuring optimal system-wide performance. Experiments demonstrate that the dense knowledge rewards generated by the CRA efficiently guide the policy agents’ learning, leading to faster convergence compared to baseline methods. The CenRA also performs superiorly and robustly in transferring acquired knowledge to new tasks.
CenRA’s current limitation is its requirement for uniform state and action shapes across tasks. Future improvements could explore preprocessing techniques to adapt the framework for varying state and action shapes to broaden CenRA’s applicability. Additionally, the fixed trade-off between similarity weight and performance weight may not be ideal. A more flexible approach, potentially through adaptive regulation of these weights, could further enhance the framework. Besides, the performance weight might favor underperforming tasks, achieving overall balance, but potentially limiting the upper bound of top-performing tasks. This issue could necessitate a better mechanism for trade-offs.
@@ Appendix A Mutli-Task Experimental Configurations
We conduct experiments in four domains with multiple tasks: Meta-World, 2DMaze, 3DPickup, and MujocoCar. The detailed configurations of each task are illustrated in Figure 8. The Meta-World tasks illustration is adapted from (Yu et al., 2020b, ).
Figure 8: Illustration of multiple tasks in different domains in our experiments.
Appendix B Experimental Configurations